# Charge constraint on M87* with twisted light

### v1

@@ Abstract
A few years ago, a significant step forward has been made in testing general relativity at the strongest field strengths, thanks to the development of Very Long Baseline Interferometry (VLBI). This advancement has enabled the capture of horizon-scale images of supermassive black holes. As a result, a new era has been ushered in, providing new opportunities to further test fundamental physics. In particular, a major milestone in this field was achieved using the Event Horizon Telescope (EHT), a millimeter VLBI array with Earth-scale baselines [7]. In 2019, the EHT collaboration produced the first resolved images of the near-horizon region of M87* [8, 13] (see also Ref. [14], which presents the original idea underlying such observations). Subsequent observations revealed its magnetic field structure [15, 16], further characterizing the black hole’s phenomenology. More recently, the first images of Sgr A* were obtained from EHT data, providing another crucial dataset for testing gravity and fundamental physics [17, 22]. These horizon-scale images reveal key features of black holes, most notably a bright emission encircling a central brightness depression, the latter being known as the ‘shadow’ of the black hole (for detailed reviews on the black hole shadow and related concepts, see, e.g., Refs. [23, 27]).
@@
First theorized by Abraham in 1914 [43], OAM is one of the two components of total electromagnetic angular momentum, the other being spin angular momentum (SAM). Over the years, OAM has found a broad range of applications across multiple fields, including nanotechnology [44, 45], quantum cryptography [46, 47] and telecommunications [48, 56], extending beyond classical multiplexing schemes [57]. Only very recently has OAM been explored in astronomy and astrophysics, as previously suggested in Ref. [58]. Theoretical predictions and experimental verification confirm that light passing near a rotating black hole (modeled as a Kerr black hole [59]) acquires a twist in its spatial phase distribution due to the rotating spacetime, enabling the extraction of information about the black hole’s rotation [60, 62]. Moreover, exploiting the analogy between gravitational fields and optical media within the general relativistic geometric optics framework [63, 64], simple formulas have been proposed to characterize the OAM of radiation from a rotating compact object [65]. This allows us to construct a refined criterion for modeling Kerr-like geometries and, as we show here, to derive an explicit bound on the electric charge of M87*.
The paper is organized as follows. In Sec. II, some fundamental concepts regarding SAM and OAM are reviewed; in particular, the so-called spin-orbit coupling process, which occurs when light traverses anisotropic and inhomogeneous media, is introduced. In Sec. III, it is shown that this same process also takes place when light passes near rotating compact objects, ultimately leading to a simple formula for calculating OAM. In Sec. IV, starting from this formula, a constraint on the electric charge of M87* is derived. Finally, Sec. V is devoted to concluding remarks.
@@ 1 SAM-to-OAM conversion in inhomogeneous media
An incident left-circularly polarized beam described by by a Jones electric field vector $\mathbf{E}_{\text{in}}=E_{0}\times[1,i]^{T}$, emerging from the medium, transforms as [66]
@@ 2 Interplay between SAM and OAM in rotating black hole spacetimes
    A light beam tipically consists of multiple OAM modes, each associated with a distinct $m$ value. The OAM content can be characterized by the spiral spectrum, which is a histogram of the weights of each OAM component present in the received beam (see, e.g., the lower panel of Fig. 3 in Ref. [61]). The overall OAM is expressed as $\langle m\rangle$, reflecting the average based on the distribution of these modes. In the case at hand, since the OAM distribution is dominated by the modes $m=0$ and $m=1$ [61], the average OAM can be estimated by considering the ratio between the weights of these two components [62]:
@@
Figure 1: Average OAM ($\langle m\rangle$) of radiation emitted near a charged rotating black hole as a function of the radial coordinate $r$ measured from the black hole center (in units of $3M$). Different curves correspond to varying electric charges and spin parameters. The black hole mass is normalized to $M=1$ and and the observer’s viewing angle is fixed at $\vartheta_{0}=0$, aligned with the black hole’s spin axis (polar viewing angle). The graph illustrates how the OAM content of the emitted light changes with both charge and spin, highlighting its sensitivity to these parameters. Three key features emerge clearly: (i) the OAM increases as the spin parameter $a$ increases, (ii) the OAM increases as the radial coordinate $\bar{r}$ decreases (approaching the black hole), and (iii) the OAM increases as the electric charge $\mathcal{Q}$ increases.

### v2

@@ Abstract
A few years ago, a significant step forward has been made in testing general relativity at the strongest field strengths, thanks to the development of Very Long Baseline Interferometry (VLBI). This advancement has enabled the capture of horizon-scale images of supermassive black holes. As a result, a new era has been ushered in, providing new opportunities to further test fundamental physics. In particular, a major milestone in this field was achieved using the Event Horizon Telescope (EHT), a millimeter VLBI array with Earth-scale baselines [7]. In 2019, the EHT collaboration produced the first resolved images of the near-horizon region of M87* [8][13] (see also Ref. [14], which presents the original idea underlying such observations). Subsequent observations revealed its magnetic field structure [15, 16], further characterizing the black hole’s phenomenology. More recently, the first images of Sgr A* were obtained from EHT data, providing another crucial dataset for testing gravity and fundamental physics [17][22]. These horizon-scale images reveal key features of black holes, most notably a bright emission encircling a central brightness depression, the latter being known as the ‘shadow’ of the black hole (for detailed reviews on the black hole shadow and related concepts, see, e.g., Refs. [23][27]).
@@
First theorized by Abraham in 1914 [43], OAM is one of the two components of total electromagnetic angular momentum, the other being spin angular momentum (SAM). Over the years, OAM has found a broad range of applications across multiple fields, including nanotechnology [44, 45], quantum cryptography [46, 47] and telecommunications [48][56], extending beyond classical multiplexing schemes [57]. Only very recently has OAM been explored in astronomy and astrophysics, as previously suggested in Ref. [58]. Theoretical predictions and experimental verification confirm that light passing near a rotating black hole (modeled as a Kerr black hole [59]) acquires a twist in its spatial phase distribution due to the rotating spacetime, enabling the extraction of information about the black hole’s rotation [60][62]. Moreover, exploiting the analogy between gravitational fields and optical media within the general relativistic geometric optics framework [63, 64], simple formulas have been proposed to characterize the OAM of radiation from a rotating compact object [65]. This allows us to construct a refined criterion for modeling Kerr-like geometries and, as we show here, to derive an explicit bound on the electric charge of M87*.
The paper is organized as follows. In Sec. 1, some fundamental concepts regarding SAM and OAM are reviewed; in particular, the so-called spin-orbit coupling process, which occurs when light traverses anisotropic and inhomogeneous media, is introduced. In Sec. 2, it is shown that this same process also takes place when light passes near rotating compact objects, ultimately leading to a simple formula for calculating OAM. In Sec. 3, starting from this formula, a constraint on the electric charge of M87* is derived. Finally, Sec. 4 is devoted to concluding remarks.
@@ 1 SAM-to-OAM conversion in inhomogeneous media
An incident left-circularly polarized beam described by a Jones electric field vector $\mathbf{E}_{\text{in}}=E_{0}\times[1,i]^{T}$, emerging from the medium, transforms as [66]
@@ 2 Interplay between SAM and OAM in rotating black hole spacetimes
    A light beam typically consists of multiple OAM modes, each associated with a distinct $m$ value. The OAM content can be characterized by the spiral spectrum, which is a histogram of the weights of each OAM component present in the received beam (see, e.g., the lower panel of Fig. 3 in Ref. [61]). The overall OAM is expressed as $\langle m\rangle$, reflecting the average based on the distribution of these modes. In the case at hand, since the OAM distribution is dominated by the modes $m=0$ and $m=1$ [61], the average OAM can be estimated by considering the ratio between the weights of these two components [62]:
@@
Figure 1: Average OAM ($\langle m\rangle$) of radiation emitted near a charged rotating black hole as a function of the radial coordinate $r$ measured from the black hole center (in units of $3M$). Different curves correspond to varying electric charges and spin parameters. The black hole mass is normalized to $M=1$ and the observer’s viewing angle is fixed at $\vartheta_{0}=0$, aligned with the black hole’s spin axis (polar viewing angle). The graph illustrates how the OAM content of the emitted light changes with both charge and spin, highlighting its sensitivity to these parameters. Three key features emerge clearly: (i) the OAM increases as the spin parameter $a$ increases, (ii) the OAM increases as the radial coordinate $\bar{r}$ decreases (approaching the black hole), and (iii) the OAM increases as the electric charge $\mathcal{Q}$ increases.